# Quantum simulation of the nonlinear Schrödinger equation via measurement-induced potential reconstruction

### v1

@@ Abstract
The nonlinear Schrödinger equation (NLSE) is a fundamental model that describes diverse complex phenomena in nature. However, simulating the NLSE on a quantum computer is inherently challenging due to the presence of the nonlinear term. We propose a hybrid quantum-classical framework for simulating the NLSE based on the split-step Fourier method. During the linear propagation step, we apply the kinetic evolution operator to generate an intermediate quantum state. Subsequently, the Hadamard test is employed to measure the Fourier components of low-wavenumber modes, enabling the efficient reconstruction of nonlinear potentials. The phase transformation corresponding to the reconstructed potential is then implemented via a quantum circuit using the phase kickback technique. To validate the efficacy of the proposed algorithm, we numerically simulate the evolution of a Gaussian wave packet, a soliton wave, and the wake flow past a cylinder. The simulation results demonstrate excellent agreement with the corresponding classical solutions. This work provide a concrete basis for analyzing accuracy-cost trade-offs in quantum-classical simulations of nonlinear dispersive wave dynamics.
@@ I Introduction
Here, we present a hybrid quantum-classical split-step spectral framework for simulating the NLSE. The method combines quantum Fourier transform (QFT) for linear propagation with measurement-assisted local-phase updates for the nonlinear regime, thereby maintaining globally unitary evolution through sequences of quantum gates. We devise and analyze a spectral filtering strategy that retains dominant modes while reducing effective dimensionality, enabling stable temporal integration under restricted qubit and measurement constraints. Computational complexity analysis and comparisons with classical algorithms are provided. Furthermore, we validate the efficacy of the present framework by numerically simulating 1D solitons, 2D Gaussian wave packets, and 2D wake flow past a cylinder.
@@ III.3 Complexity analysis
$T\approx\sum_{t_{j}=1}^{N_{t}}M\cdot N_{\text{shots}}\cdot\mathcal{O}(t_{j}\log^{2}N)\approx\mathcal{O}\left(N_{t}^{2}M\log^{2}N/\varepsilon^{2}\right).$ (14)
@@ IV.1 Numerical setup
We validate the method using three test cases: a 1D soliton, a 2D Gaussian wave packet, and a moving 2D cylindrical obstacle within a superfluid BEC. For all cases, periodic boundary conditions are imposed to facilitate the numerical implementation.
@@
Finally, we evaluate the necessity of renormalization within the filtered-QSSFM workflow. Figure 6 compares three simulation regimes using $n=8$ and $m=4$. The standard QSSFM (see Fig. 6(a)) serves as the baseline, accurately reproducing both the amplitude and phase of the SSFM reference. In the filtered-QSSFM without normalization (see Fig. 6(b)), the nonlinear term is evaluated directly using the unnormalized retained modes. Consequently, numerical errors accumulate over successive iterations, resulting in pronounced deviations from the exact dynamics. Figure 6(c) illustrates the efficacy of the renormalization strategy: by normalizing the reconstructed wave function prior to evaluating the nonlinear operator, the simulation stability is significantly improved. Although the filtering step inherently truncates high-frequency energy, resulting in a systematic, albeit minor, amplitude decay, the renormalization process numerically compensates for this norm loss. This compensation ensures that the filtered-QSSFM maintains close agreement with the classical SSFM reference, thereby avoiding the divergence observed in the unnormalized approach.
@@
In this 2D case, the quantum register comprises $n=14$ qubits, distributed as $n_{x}=7$ and $n_{y}=7$ for the respective spatial coordinates. For the filtering stage, the number of retained qubits is set to $m=6$, corresponding to $m_{x}=3$ and $m_{y}=3$ active spectral modes per dimension. Figure 7(a) displays the time evolution obtained via the standard QSSFM, while Fig. 7(b) presents the results generated by the classically simulated filtered-QSSFM. The corresponding error distribution is plotted in Fig. 7(c), where the deviation is quantified as
@@
IV.4 2D wake flow past a cylinder
@@
Figure 8 presents a comparison between the classical SSFM reference solutions and the results obtained via the filtered-QSSFM. The density profile $|\psi|^{2}$, relative phase distribution, and vorticity field are illustrated in Figs. 8(a), 8(b), and 8(c), respectively. Distinct vortex shedding is observed, demonstrating that the proposed framework effectively captures the characteristic vortex generation and shedding dynamics inherent to the 2D BEC flow past an obstacle. These results indicate that the present methodology is extensible to higher spatial dimensions.

### v2

# Quantum computing of the nonlinear Schrödinger equation via measurement-induced potential reconstruction

@@ Abstract
The nonlinear Schrödinger equation (NLSE) is a fundamental model that describes diverse complex phenomena in nature. However, simulating the NLSE on a quantum computer is inherently challenging due to the presence of the nonlinear term. We propose a hybrid quantum-classical framework for simulating the NLSE based on the split-step Fourier method. During the linear propagation step, we apply the kinetic evolution operator to generate an intermediate quantum state. Subsequently, the Hadamard test is employed to measure the Fourier components of low-wavenumber modes, enabling the efficient reconstruction of nonlinear potentials. The phase transformation corresponding to the reconstructed potential is then implemented via a quantum circuit using the phase kickback technique. To validate the efficacy of the proposed algorithm, we numerically simulate the evolution of a Gaussian wave packet, a soliton wave, snake instability, and the wake flow past a cylinder. The simulation results demonstrate excellent agreement with the corresponding classical solutions. This work provide a concrete basis for analyzing accuracy-cost trade-offs in quantum-classical simulations of nonlinear dispersive wave dynamics.
@@ I Introduction
Here, we present a hybrid quantum-classical split-step spectral framework for simulating the NLSE. The method combines quantum Fourier transform (QFT) for linear propagation with measurement-assisted local-phase updates for the nonlinear regime, thereby maintaining globally unitary evolution through sequences of quantum gates. We devise and analyze a spectral filtering strategy that retains dominant modes while reducing effective dimensionality, enabling stable temporal integration under restricted qubit and measurement constraints. Computational complexity analysis and comparisons with classical algorithms are provided. Furthermore, we validate the efficacy of the present framework by numerically simulating 1D solitons, 2D Gaussian wave packets, 2D snake instability, and 2D wake flow past a cylinder.
@@ III.3 Complexity analysis
$T\approx\sum_{t_{j}=1}^{N_{t}}MN_{\text{shots}}\mathcal{O}(t_{j}\log^{2}N)\approx\mathcal{O}\left(N_{t}^{2}M\log^{2}N/\varepsilon^{2}\right).$ (14)
@@ IV.1 Numerical setup
We validate the method using four test cases: a 1D soliton, a 2D Gaussian wave packet, 2D snake instability and a moving 2D cylindrical obstacle within a superfluid BEC. For all cases, periodic boundary conditions are imposed to facilitate the numerical implementation.
@@
Finally, we evaluate the necessity of renormalization within the filtered-QSSFM workflow. Figure 6 compares three simulation regimes using $n=8$ and $m=4$. The standard QSSFM (see Fig. 6(a)) serves as the baseline, accurately reproducing both the amplitude and phase of the SSFM reference. In the filtered-QSSFM without normalization (see Fig. 6(b)), the nonlinear term is evaluated directly using the unnormalized retained modes. Consequently, numerical errors accumulate over successive iterations, resulting in pronounced deviations from the exact dynamics. Figure 6(c) illustrates the efficacy of the renormalization strategy, where the simulation stability is significantly improved by normalizing the reconstructed wave function prior to evaluating the nonlinear operator. Although the filtering step inherently truncates high-frequency energy, resulting in a systematic, albeit minor, amplitude decay, the renormalization process numerically compensates for this norm loss. This compensation ensures that the filtered-QSSFM maintains close agreement with the classical SSFM reference, thereby avoiding the divergence observed in the unnormalized approach.
@@
Figure 8: Simulation results for the 2D snake instability evolution at $t=0,50$, and 6060 using $n_{x}=n_{y}=8$ qubits: (a) the QSSFM, (b) the filtered-QSSFM with $m_{x}=m_{y}=5$.
@@
In this 2D case, the quantum register comprises $n=14$ qubits, distributed as $n_{x}=n_{y}=7$ for the respective spatial coordinates. For the filtering stage, the number of retained qubits is set to $m=6$, corresponding to $m_{x}=m_{y}=3$ active spectral modes per dimension. Figure 7(a) displays the time evolution obtained via the standard QSSFM, while Fig. 7(b) presents the results generated by the classically simulated filtered-QSSFM. The corresponding error distribution is plotted in Fig. 7(c), where the deviation is quantified as
@@
IV.4 2D snake instability
We investigate the snake instability that emerges from the evolution of a dark soliton solution of the NLSE with parameters $g=1$ and $V=1$. The computational domain is taken as $x,y\in[-25,25]$, and the time step is $\Delta t=0.01$. We use $n_{x}=n_{y}=8$ qubits for amplitude encoding, and set $m_{x}=m_{y}=5$ retained qubits in the filtered-QSSFM. The initial condition is a perturbed dark soliton
[…]
Figure 8 illustrates the results obtained from the QSSFM and filtered-QSSFM at $t=0,50$, and $60$, respectively. The initial soliton configuration is depicted at $t=0$. At $t=50$, the onset of instability becomes apparent through the pronounced transverse bending of the dark soliton, which remains spatially continuous. Subsequently, the soliton undergoes fragmentation at $t=60$, signifying the full development of the snake instability.
However, a discernible discrepancy between the results in Figs. 8(a) and (b) emerges at $t=60$. This difference is primarily attributed to high-wavenumber modes generated during the evolution of the instability. These components manifest as high-frequency vertical structures in the snapshots at $t=50$ and $t=60$. Within the filtered QSSFM framework, such high-frequency contributions are truncated during the mode-filtering stage. Consequently, the cumulative removal of these modes leads to increasingly prominent deviations between the two methods as the instability develops.
IV.5 2D wake flow past a cylinder
@@
Figure 9 presents a comparison between the classical SSFM reference solutions and the results obtained via the filtered-QSSFM. The density profile $|\psi|^{2}$, relative phase distribution, and vorticity field are illustrated in Figs. 9(a), 9(b), and 9(c), respectively. Distinct vortex shedding is observed, demonstrating that the proposed framework effectively captures the characteristic vortex generation and shedding dynamics inherent to the 2D BEC flow past an obstacle. These results indicate that the present methodology is extensible to higher spatial dimensions.